A square pool has a side length of X feet. A uniform border around the pool is 1 foot wide. The total area of the pool and the border is 361 square feet. What is the area of the pool?
Answer by mananth(16946)

A square pool has a side length of X feet.
A uniform border around the pool is 1 foot wide.
length including the border = (x+2) feet
Area = (x+2)(x+2)
(x+2)^2 = 361 ft^2
taking square root
x+2 = +/- sqrt(361)
x+2=+/-19
x= 17 taking positive value
Area of pool = 17^2 = 289 ft^2